OpenStudy (anonymous):

Solve please! \[\left| 4x-5 \right|=10\]

OpenStudy (mathmate):

case 1: 4x-5=10 => x=15/4 case 2: -(4x-5)=10 => x=-5/4

OpenStudy (anonymous):

Nice we are going good! :D

OpenStudy (mathmate):

Great, but I've got to hit the sack now. Talk to you soon!

OpenStudy (anonymous):

ok man thank you very much. Sweet dreams (no homo) haha